# Dynamical arrest in active nematic turbulence

### v2

@@ Abstract
Active fluids display spontaneous turbulent-like flows known as active turbulence. Recent work revealed that these flows have universal features, independent of the material properties and of the presence of topological defects. However, the differences between defect-laden and defect-free active turbulence remain largely unexplored. Here, by means of large-scale numerical simulations, we show that defect-free active nematic turbulence can undergo dynamical arrest. We find that flow alignment—the tendency of nematics to reorient under shear—enhances large-scale jets in contractile rodlike systems while promoting arrested flow patterns in extensile systems. Our results reveal a mechanism of labyrinthine pattern formation produced by an emergent topology of nematic domain walls that partially suppresses chaotic flows. Taken together, our findings call for the experimental realization of defect-free active nematics, and suggest that topological defects enable turbulence by preventing dynamical arrest.
@@
Active turbulence encompasses systems with different symmetries, including polar or nematic orientational order, each exhibiting unique statistical properties 1. Recent work showed that the velocity power spectrum of active nematic turbulence features scaling laws with universal exponents independent of the material properties of the fluid such as its viscosity and activity 32; 33; 22. Whereas most experiments and simulations prominently feature topological defects 34, turbulence with the same universal exponents was also found for strongly-ordered nematics in the absence of defects 33. However, how the material properties and the presence or absence of defects affect other, non-universal features of the flow field remains unclear.
@@
We identify two distinct regimes depending on the parameter combination $\text{sign}(\zeta)\nu$, where $\zeta$ is the active stress parameter, with $\zeta>0$ for extensile and $\zeta<0$ for contractile stresses, respectively. We discuss the case of rod-like nematic components, for which $\nu\leq 0$. In the flow-aligning regime, where $\nu<-1$, we find that contractile nematics (with $\zeta\nu>0$) exhibit strongly chaotic flows and nematic distortions, which largely decorrelate from one another. In this regime, nematic domain walls—lines of strong variation of the director—appear, split, and dissolve, forming a dynamically reorganizing pattern (Figs. 1 to 1, Supplementary Movie 1). In contrast, for extensile aligning nematics ($\zeta\nu<0$), we find that domain walls are strongly stabilized by the flow they entrain. The walls typically grow and branch but neither break nor merge. As walls avoid each other, they become arrested (or grid-locked) in a space-filling tree-like pattern (Figs. 1 to 1, Supplementary Movie 2). Once arrested, the pattern exhibits slow residual dynamics that shares some features with ageing phenomena in glassy systems. The special case $\nu=0$ has been the focus of previous work33 and is considered here as a reference. The dynamics in this case lie between these two extremes while also manifesting comparatively shorter wavelengths (Supplementary Fig. 1, Supplementary Movie 3).
We uncover that the dynamical arrest stems from the emergence of an effective topology, characterized by a set of connectivity rules for the domain walls imposed by the combined effects of activity and flow alignment. In particular, we demonstrate that, in the absence of topological defects, the arrested state is organized by local nematic structures, which we name pseudo-defects. We discuss their topological interpretation and show that they coexist with actual defects in experiments on microtubule-based active nematics. While future studies could clarify which features of defect-free active flows may also be present in active nematics with defects, achieving defect-free active nematics remains a compelling experimental challenge.
Hydrodynamics of defect-free active nematics
We generalize our previous minimal model of active nematic turbulence 33 by adding flow alignment and the Ericksen stress 35. We work at vanishing Reynolds number, for which momentum balance reads:
@@
The pressure $P$ enforces the incompressibility condition $\partial_{\alpha}v_{\alpha}=0$ of the flow field $\bm{v}$. The next two terms are the divergence of the antisymmetric part of the stress tensor, and of the symmetric Ericksen stress, respectively35; 36:
@@
Here, we assumed that the fluid is deep in the nematic phase so that the director field has a fixed modulus $|\bm{n}|=1$. This constraint is imposed by the second term through the Lagrange multiplier $h_{\parallel}^{0}$. We start with a continuous director field, which precludes the generation of topological defects 33; 37.
The last term in Eq. 1 arises from the symmetric part of the deviatoric stress, which is given by the constitutive equation 38; 39; 40; 36
@@
Nondimensionalizing the equations, we rescale length by the system size $L$, time by the active time $\tau_{\text{a}}=\eta/|\zeta|$, pressure by the active stress $|\zeta|$, and orientational field by $K/L^{2}$. To eliminate the pressure, we take the curl of Eq. 1 and obtain a Poisson equation for the vorticity $\omega$, which we write in terms of the stream function $\psi$ (defined by $v_{x}=\partial_{y}\psi$ and $v_{y}=-\partial_{x}\psi$) 37, as $\omega=-\nabla^{2}\psi$:
$\nabla^{2}\omega=-\nabla^{4}\psi=-S\partial_{\alpha}\partial_{\beta}\bar{q}_{\alpha\beta}+R\nu\partial_{\alpha}\partial_{\beta}(\bar{q}_{\alpha\beta}h_{\parallel}-\hat{q}_{\alpha\beta}h_{\perp})\\ +\frac{R}{2}\nabla^{2}h_{\perp}+R\epsilon_{\alpha\beta}(\partial_{\alpha}h_{\parallel})\partial_{\beta}\theta.$ (6)
Here, $S$ is the sign of the active stresses, with $S=\pm 1$ for extensile and contractile stresses, respectively, and $\theta$ is the nematic director angle: $\bm{n}=(\cos\theta,\,\sin\theta)$. The right-hand side contains four sources of vorticity, arising respectively from active, flow-alignment, antisymmetric, and Ericksen stresses. We used the totally antisymmetric tensor $\epsilon_{\alpha\beta}$ to define $\bar{q}_{\alpha\beta}\equiv-\epsilon_{\alpha\gamma}\hat{q}_{\gamma\beta}$. We also introduced the parallel and perpendicular components of the orientational field 37:
@@
Finally, the director dynamics Eq. 5 breaks down into a component parallel to $\bm{n}$, which specifies $h_{\parallel}$ as given in Eq. 7, and a component perpendicular to $\bm{n}$, which determines the dynamics of the nematic angle $\theta$ 37:
@@
We numerically integrate Eqs. 6 to 8 in a square with periodic boundary conditions using a hybrid numerical scheme that combines a pseudo-spectral method for Eq. 6 with the finite element method for Eq. 8 (Methods .1). Unless otherwise specified, we set $R=1$ and $A=3.2\times 10^{5}$, for which the system was shown to develop large-scale active turbulence when $\nu=0$ 33. The dynamics are invariant under the transformation ($\nu\rightarrow-\nu$, $S\rightarrow-S$, $\theta\rightarrow\theta\pm\pi/2$), meaning that changing the signs of $\nu$ and $S$ simultaneously leaves the system invariant up to a rotation 41; 42; 37.
@@
Figure 1: Strong and arrested regimes of active nematic turbulence. 11, Snapshots from simulations of defect-free active nematic turbulence in contractile (11) and extensile (11) flow-aligning systems. Parameter values were set to $R=1$, $\nu=-1.1$, and $A=3.2\times 10^{5}$. Top panels (1,1) show the flow field; black curves are streamlines, and the color indicates the speed (see Supplementary Movies 1 and 2). Middle panels (1,1) show the Frank free energy density $\sim|\nabla\theta|^{2}$, with high-intensity lines corresponding to nematic domain walls (see Supplementary Movies 1 and 2). Bottom panels (1,1) are zooms highlighting the type of nematic distortion as well as the interplay between nematic walls and flows. The gray-scale background is the line integral convolution representation of the director field $\bm{n}$. Magenta and cyan intensities respectively represent splay $(\nabla\cdot\bm{n})^{2}$ and bend |×$|\nabla\times\bm{n}|^{2}$ contributions to the Frank energy density. The black arrows represent the flow field $\bm{v}$, which localizes along the nematic walls in the arrested regime. Black circles indicate stagnation points of the flow. White scale bar represents the selected wavelength $\lambda_{i}$. 11, Spectra characterizing fully-developed active nematic turbulence (see details in Methods .2). The lines in (1,1) represent a smoothed (Gaussian) interpolation of the computed data points. We compare the flow-aligning contractile (red, as in 11 and Supplementary Movie 1) and extensile (blue, as in 11 and Supplementary Movie 2) cases with the $\nu=0$ case (black, as in Supplementary Fig. 1 and Supplementary Movie 3), for which contractile and extensile stresses are equivalent up to a rotation41; 42; 37 . 1, Velocity power spectrum on a log-log scale, showing (i) the universal low-$q$ scaling law and (ii) the distinct organization of flows across scales in the different cases. The wider scaling regime in the contractile case captures the strong large-scale jets (see 1). The peak in the extensile case is representative of wall streams (see 1). 1, Frank energy spectrum, showing that (i) the selected wavelength (peak position) depends on $\nu$ but not on the sign of active stress, and (ii) the peak width depends on the sign of active stress when $\nu\neq 0$. 1, Spectrum of correlation times associated with the flow $\bm{v}$ (light colored points and lines) and the nematic tensor $\hat{q}_{\alpha\beta}$ (darker points and lines). This log-log plot reveals strong differences in decay times between the regimes, as well as the differences between the flow and nematic tensor within a regime. Correlation times are extracted from exponential fits to the corresponding space-time autocorrelation functions in Fourier space (see Methods .2).
Strong large-scale turbulence
@@
We first study the case $S\nu>1$, realized with contractile flow-aligning rods ($S=-1$ and $\nu=-1.1$). We find a chaotic flow field with vortices at many scales and strong (albeit transient) large-scale jets (Fig. 1). These flow patterns stem from long-range hydrodynamic interactions and are largely uncorrelated with the director field (Fig. 1), which also evolves chaotically. The Frank energy density concentrates along domain walls (Figs. 1 and 1)—well-known regions of strong orientation gradients 43; 44; 45. Such walls feature a typical separation length, corresponding to the peak position in the Frank energy spectrum (Fig. 1, red), coinciding with the peak of the nematic tensor power spectrum (Supplementary Fig. 2, red). Compared with the $\nu=0$ case33, the nematic pattern in the $S\nu>1$ regime appears more broken, with fragmented walls coexisting with wall-free regions (dark patches in Fig. 1). As a consequence, the total Frank free energy is smaller, and its peak is shifted to larger scales (Fig. 1, red). A dynamically-rearranging pattern emerges as walls spontaneously appear, split, and dissolve (Supplementary Movies 1 and 4, left). Both the flows and the nematic tensor are very dynamic, as confirmed by their short correlation time at all length scales (Fig. 1, red).
@@
In recent work we have shown that, for $S\nu>1$, both the quiescent state with uniform orientation and the flowing state with splaying domain walls are non-linearly unstable against one-dimensional modulations37. Thus, strong enough perturbations can trigger sudden transitions between these states. For this reason, we assert that domain wall patterns can be disrupted by the strong turbulent flows. In the two-dimensional simulations shown here, we observe the coexistence of uniform and wall-laden patches that continuously rearrange. We conclude that the metastibility associated with this regime strongly enhances the chaotic flows, particularly at large-scales, as indicated by the wide $q^{-1}$ scaling regime in Fig. 1, red.
Arrested turbulence
@@
In contrast to the non-arrested case, here a connected wall pattern extends across the entire system (Fig. 1). The Frank free energy spectrum is more narrowly peaked at the characteristic length scale (Fig. 1, blue), compared to the non-arrested case (Fig. 1, red). The same peak is also prominent in the nematic tensor power spectrum (Supplementary Fig. 2, blue). Strikingly, the nematic wall pattern becomes far less dynamic; the walls and their associated streams become locked in a tree-like pattern. We term this phenomenon dynamical arrest. The weak large-scale chaotic flows make the pattern fluctuate only slightly (Supplementary Movies 2 and 4, right). The dynamical arrest manifests in much longer correlation times, especially of the nematic tensor field at large scales (Fig. 1, blue). The tree-like pattern of domain walls is characterized by unique connectivity properties, which we elaborate on later. We conclude that, in the absence of topological defects, the combined effects of extensile stress and rod-like flow-alignment give rise to the arrest of active nematic turbulence.
@@
Wall dynamics and mechanism of arrest
To understand the mechanism of arrest, we note that the spontaneous-flow instability, which ultimately drives turbulence, generates shear flow 46. This shear produces antiparallel flows that concentrate along the nematic walls46; 41; 42; 47; 33; 37. The walls are directed lines, with their direction defined by the vector $(\partial_{y}\theta,-\partial_{x}\theta)$, which is perpendicular to the gradient of the angle and coincides with the flow direction in the arrested regime.
Quasi one-dimensional wall patterns tend to coarsen, leading to larger domains (Fig. 2 and Supplementary Movie 4, right). Since the director orientation is roughly uniform within such domains, the same instability prompts distortions in a perpendicular direction33; 47. This process induces sequential ‘zig-zag’ instabilities, leading to wall branching and progressively smaller domains, until a characteristic wavelength is selected (Fig. 2 and Supplementary Movie 4, right) 19; 33. When an open end of a branch approaches another wall perpendicularly, it is deflected and becomes antiparallel to that wall. The growth of branches typically stops when they are blocked by a preexisting branch point. Ultimately, these dynamics lead to a grid-locked pattern (Fig. 2 and Supplementary Movie 4, right).
When a growing wall is trapped by a branch point, the outflowing stream at the end point of the wall is diverted to both sides so that material can join the antiparallel streams of neighbouring walls (see Fig. 1, around the black circles). In other words, a stagnation point is formed as flows arrive in the direction of the growing wall and leave in the perpendicular direction, to which rods tends to align. For bending walls—induced by an extensile stress—the director is perpendicular to the wall and therefore aligned with the extensional axis of the flow. We thus postulate that the structure around the stagnation point acts as a trap, and its stability is key to understanding how flow alignment promotes the arrest of active turbulence.
@@
Figure 3: Domain wall nodes and their pseudo-topology. In all diagrams, black directed lines represent nematic bending walls, with gray lines tracing the director and cyan indicating the bending energy. 33, Nodes are either startpoints (green, negative pseudo-defects), branchpoints (blue, negative pseudo-defects), or endpoints (red, positive pseudo-defects). The orange curves illustrate the paths excluding walls used to define the pseudo-charge (Supplementary Note). 33, Two ways in which a startpoint-endpoint pair of opposite charge can be born. The inverse processes of (3) (complete wall dissolution) and (3) (local wall completion) result in the annihilation of such pairs. 3, Wall branching gives birth to a branchpoint-endpoint pair, also of opposite charge. A connected pair may annihilate via branch retraction. 3, A branchpoint (3) transitions into a startpoint when one outgoing wall disjoins. The inverse process corresponds to a startpoint (3) joining with a bare wall. 3, Connected branchpoints can shrink their connecting branch, transiently creating a $-2Q$ pseudo-charged structure, exchange outgoing walls and extend again in the perpendicular direction.
The nodes of nematic domain walls define pseudo-defects

### v3

@@ Abstract
Active fluids display spontaneous turbulent-like flows known as active turbulence. Recent work revealed that these flows have universal features, independent of the material properties and of the presence of topological defects. However, the differences between defect-laden and defect-free active turbulence remain largely unexplored. Here, by means of large-scale numerical simulations, we show that defect-free active nematic turbulence can undergo dynamical arrest. This state is characterized by an emergent network of nematic domain walls that channels coherent streams and suppresses chaotic flows. As the system evolves, the branched wall network produces a large-scale pattern with tree-like topological properties. We find that flow alignment—the tendency of nematics to reorient under shear—enhances large-scale chaotic jets in contractile rodlike systems while promoting dynamical arrest in extensile systems. We further show that dynamical arrest persists regardless of whether defects are prohibited by construction or simply fail to form due to a high energy cost of defect cores. Taken together, our findings reveal a striking pattern-formation mechanism, with labyrinths emerging from active turbulence, and illuminate the rich transitional regime between defect-free and defect-laden dynamics. These behaviors call for the experimental realization of active nematics at vanishing or low defect densities, and underscore that, in extensile rodlike nematics, topological defects enable turbulence by preventing dynamical arrest.
I Introduction
@@
Active turbulence encompasses systems with different symmetries, including polar or nematic orientational order, each exhibiting unique statistical properties 1. Recent work showed that the velocity power spectrum of active nematic turbulence features scaling laws with universal exponents independent of the material properties of the fluid such as its viscosity and activity 32; 33; 22. In particular, the $q^{-1}$ scaling of the velocity power spectrum at small wavenumbers $q$ was shown to be highly robust in simulations, provided the Reynolds number remains low 34. Whereas the 1-1 exponent was first predicted for defect-laden dynamics 32, turbulence with the same universal exponent was also found for strongly-ordered nematics in the absence of defects 33. However, how the material properties and the presence or absence of defects affect other, non-universal features of the flow field remains unclear.
@@
In the absence of defects, we identify two distinct regimes depending on the parameter combination $\text{sign}(\zeta)\nu$, where $\zeta$ is the active stress parameter, with $\zeta>0$ for extensile and $\zeta<0$ for contractile stresses, respectively. We discuss the case of rod-like nematic components, for which $\nu\leq 0$. In the flow-aligning regime, where $\nu<-1$, we find that contractile nematics (with $\zeta\nu>0$) exhibit strongly chaotic flows and nematic distortions, which largely decorrelate from one another. In this regime, nematic domain walls—lines of strong variation of the director—appear, split, and dissolve, forming a dynamically reorganizing pattern (Figs. 1 to 1, Movie S1). In contrast, for extensile aligning nematics ($\zeta\nu<0$), we find that domain walls are strongly stabilized by the flow they entrain. The walls typically grow and branch but neither break nor merge. As walls avoid each other, they become arrested (or grid-locked) in a space-filling tree-like pattern (Figs. 1 to 1, Movie S2). Once arrested, the pattern exhibits slow residual dynamics that shares some features with ageing phenomena in glassy systems. The special case $\nu=0$ has been the focus of previous work33 and is considered here as a reference. The dynamics in this case lie between these two extremes while also manifesting comparatively shorter wavelengths (Fig. S135, Movie S3).
We uncover that the dynamical arrest stems from the emergence of an effective topology, characterized by a set of connectivity rules for the domain walls imposed by the combined effects of activity and flow alignment. In particular, we demonstrate that, in the absence of topological defects, the arrested state is organized by local nematic structures, which we name pseudo-defects. We discuss their topological interpretation and demonstrate that they coexist with actual defects in experiments on microtubule-based active nematics.
Finally, we show that dynamical arrest is not exclusive to our minimal defect-free model, but rather a robust phenomenon that also arises in the full Q-tensor formulation of active nematics. In this more general framework, an additional control parameter—the defect core size—governs whether topological defects are energetically suppressed or favored. By tuning this parameter, we uncover a transition in which defect nucleation ramps up sharply near a threshold. As defects proliferate, the flow crosses over from a dynamically arrested state, organized into streams that are channeled through the wall network, to the more familiar regime of chaotic, vortex-dominated flows. These simulations further highlight the role of defects in enabling active turbulence and reveal a rich transitional regime where defects and domain walls can coexist.
II Director-based defect-free model
We generalize our previous minimal model of active nematic turbulence 33 by adding flow alignment and the Ericksen stress 36. This yields the active Ericksen-Leslie liquid crystal model, widely regarded as a paradigm in this field1.
We work at vanishing Reynolds number, for which momentum balance reads:
@@ II Director-based defect-free model
The pressure $P$ enforces the incompressibility condition $\partial_{\alpha}v_{\alpha}=0$ of the flow field $\bm{v}$. The next two terms are the divergence of the antisymmetric part of the stress tensor, and of the symmetric Ericksen stress, respectively36; 37:
@@
Here, we assumed that the fluid is deep in the nematic phase so that the director field has a fixed modulus $|\bm{n}|=1$. This constraint is imposed by the second term through the Lagrange multiplier $h_{\parallel}^{0}$. We start with a continuous director field, which precludes the generation of topological defects 33; 38. Note that this constraint is lifted in our analysis of the full Q-tensor model, presented in the final section of this paper.
The last term in Eq. 1 arises from the symmetric part of the deviatoric stress, which is given by the constitutive equation 39; 40; 41; 37
@@
Nondimensionalizing the equations, we rescale length by the system size $L$, time by the active time $\tau_{\text{a}}=\eta/|\zeta|$, pressure by the active stress $|\zeta|$, and orientational field by $K/L^{2}$. To eliminate the pressure, we take the curl of Eq. 1 and obtain a Poisson equation for the vorticity $\omega$, which we write in terms of the stream function $\psi$ (defined by $v_{x}=\partial_{y}\psi$ and $v_{y}=-\partial_{x}\psi$) 38, as $\omega=-\nabla^{2}\psi$:
$\nabla^{2}\omega=-\nabla^{4}\psi=-S\partial_{\alpha}\partial_{\beta}\bar{q}_{\alpha\beta}+R\nu\partial_{\alpha}\partial_{\beta}(\bar{q}_{\alpha\beta}h_{\parallel}-\hat{q}_{\alpha\beta}h_{\perp})\\ +\frac{R}{2}\nabla^{2}h_{\perp}+R\epsilon_{\alpha\beta}(\partial_{\alpha}h_{\perp})\partial_{\beta}\theta.$ (6)
Here, $S$ is the sign of the active stresses, with $S=\pm 1$ for extensile and contractile stresses, respectively, and $\theta$ is the nematic director angle: $\bm{n}=(\cos\theta,\,\sin\theta)$. The right-hand side contains four sources of vorticity, arising respectively from active, flow-alignment, antisymmetric, and Ericksen stresses. We used the totally antisymmetric tensor $\epsilon_{\alpha\beta}$ to define $\bar{q}_{\alpha\beta}\equiv-\epsilon_{\alpha\gamma}\hat{q}_{\gamma\beta}$. We also introduced the parallel and perpendicular components of the orientational field 38:
@@
Finally, the director dynamics Eq. 5 breaks down into a component parallel to $\bm{n}$, which specifies $h_{\parallel}$ as given in Eq. 7, and a component perpendicular to $\bm{n}$, which determines the dynamics of the nematic angle $\theta$ 38:
@@
We numerically integrate Eqs. 6 to 8 in a square with periodic boundary conditions using a hybrid numerical scheme that combines a pseudo-spectral method for Eq. 6 with the finite element method for Eq. 8 (Appendix A). Unless otherwise specified, we set $R=1$ and $A=3.2\times 10^{5}$, for which the system was shown to develop large-scale active turbulence when $\nu=0$ 33. The dynamics are invariant under the transformation ($\nu\rightarrow-\nu$, $S\rightarrow-S$, $\theta\rightarrow\theta\pm\pi/2$), meaning that changing the signs of $\nu$ and $S$ simultaneously leaves the system invariant up to a rotation 42; 43; 38.
@@
Figure 1: Strong and arrested regimes of active nematic turbulence. 11, Snapshots from simulations of defect-free active nematic turbulence in contractile (11) and extensile (11) flow-aligning systems. Parameter values were set to $R=1$, $\nu=-1.1$, and $A=3.2\times 10^{5}$. Top panels (1,1) show the flow field; black curves are streamlines, and the color indicates the speed (see Movies S1 and S2). Middle panels (1,1) show the Frank free energy density $\sim|\nabla\theta|^{2}$, with high-intensity lines corresponding to nematic domain walls (see Movies S1 and S2). Bottom panels (1,1) are zooms highlighting the type of nematic distortion as well as the interplay between nematic walls and flows. The gray-scale background is the line integral convolution representation of the director field $\bm{n}$. Magenta and cyan intensities respectively represent splay $(\nabla\cdot\bm{n})^{2}$ and bend $|\nabla\times\bm{n}|^{2}$ contributions to the Frank energy density. The black arrows represent the flow field $\bm{v}$, which localizes along the nematic walls in the arrested regime. Black circles indicate stagnation points of the flow. White scale bar represents the selected wavelength $\lambda_{i}$. 11, Spectra characterizing fully-developed active nematic turbulence (see details in Appendix B). The lines in (1,1) represent a smoothed (Gaussian) interpolation of the computed data points. We compare the flow-aligning contractile (red, as in 11 and Movie S1) and extensile (blue, as in 11 and Movie S2) cases with the $\nu=0$ case (black, as in Fig. S135 and Movie S3), for which contractile and extensile stresses are equivalent up to a rotation42; 43; 38 . 1, Velocity power spectrum on a log-log scale, showing (i) the universal low-$q$ scaling law and (ii) the distinct organization of flows across scales in the different cases. The wider scaling regime in the contractile case captures the strong large-scale jets (see 1). The peak in the extensile case is representative of wall streams (see 1). 1, Frank energy spectrum, showing that (i) the selected wavelength (peak position) depends on $\nu$ but not on the sign of active stress, and (ii) the peak width depends on the sign of active stress when $\nu\neq 0$. 1, Spectrum of correlation times associated with the flow $\bm{v}$ (light colored points and lines) and the nematic tensor $\hat{q}_{\alpha\beta}$ (darker points and lines). This log-log plot reveals strong differences in decay times between the regimes, as well as the differences between the flow and nematic tensor within a regime. Correlation times are extracted from exponential fits to the corresponding space-time autocorrelation functions in Fourier space (see Appendix B).
III Regimes of defect-free active turbulence
III.1 Strong large-scale turbulence
@@
We first study the case $S\nu>1$, realized with contractile flow-aligning rods (S=1S=-1 and $\nu=-1.1$). We find a chaotic flow field with vortices at many scales and strong (albeit transient) large-scale jets (Fig. 1). These flow patterns stem from long-range hydrodynamic interactions and appear to be locally uncorrelated with the distortion pattern of the director field (Fig. 1), which also evolves chaotically. The Frank energy density concentrates along domain walls (Figs. 1 and 1)—well-known regions of strong orientation gradients 44; 45; 46. Such walls feature a typical separation length, corresponding to the peak position in the Frank energy spectrum (Fig. 1, red), coinciding with the peak of the nematic tensor power spectrum (Fig. 9, red). Compared with the $\nu=0$ case33, the nematic pattern in the $S\nu>1$ regime appears more broken, with fragmented walls coexisting with wall-free regions (dark patches in Fig. 1). As a consequence, the total Frank free energy is smaller, and its peak is shifted to larger scales (Fig. 1, red). A dynamically-rearranging pattern emerges as walls spontaneously appear, split, and dissolve (Movies S1 and S4, left). Both the flows and the nematic tensor are very dynamic, as confirmed by their short correlation time at all length scales (Fig. 1, red).
@@
In recent work we have shown that, for $S\nu>1$, both the quiescent state with uniform orientation and the flowing state with splaying domain walls are non-linearly unstable against one-dimensional modulations38. Thus, strong enough perturbations can trigger sudden transitions between these states. For this reason, we assert that domain wall patterns can be disrupted by the strong turbulent flows. In the two-dimensional simulations shown here, we observe the coexistence of uniform and wall-laden patches that continuously rearrange. We conclude that the metastibility associated with this regime strongly enhances the chaotic flows, particularly at large-scales, as indicated by the wide $q^{-1}$ scaling regime in Fig. 1, red.
III.2 Arrested turbulence
@@
In contrast to the non-arrested case, here a connected wall pattern extends across the entire system (Fig. 1). The Frank free energy spectrum is more narrowly peaked at the characteristic length scale (Fig. 1, blue), compared to the non-arrested case (Fig. 1, red). The same peak is also prominent in the nematic tensor power spectrum (Fig. 9, blue). Strikingly, the nematic wall pattern becomes far less dynamic; the walls and their associated streams become locked in a tree-like pattern. We term this phenomenon dynamical arrest. The weak large-scale chaotic flows make the pattern fluctuate only slightly (Movies S2 and S4, right). The dynamical arrest manifests in much longer correlation times, especially of the nematic tensor field at large scales (Fig. 1, blue). The tree-like pattern of domain walls is characterized by unique connectivity properties, which we elaborate on later. We conclude that, in the absence of topological defects, the combined effects of extensile stress and rod-like flow-alignment give rise to the arrest of active nematic turbulence.
@@
To understand the mechanism of arrest, we note that the spontaneous-flow instability, which ultimately drives turbulence, generates shear flow 47. This shear produces antiparallel flows that concentrate along the nematic walls47; 42; 43; 48; 33; 38. The walls are directed lines, with their direction defined by the vector $(\partial_{y}\theta,-\partial_{x}\theta)$, which is perpendicular to the gradient of the angle and coincides with the flow direction in the arrested regime.
Quasi one-dimensional wall patterns tend to coarsen, leading to larger domains (Fig. 2 and Movie S4, right). Since the director orientation is roughly uniform within such domains, the same instability prompts distortions in a perpendicular direction33; 48. This process induces sequential ‘zig-zag’ instabilities, leading to wall branching and progressively smaller domains, until a characteristic wavelength is selected (Fig. 2 and Movie S4, right) 19; 33. When an open end of a branch approaches another wall perpendicularly, it is deflected and becomes antiparallel to that wall. The growth of branches typically stops when they are blocked by a preexisting branch point. Ultimately, these dynamics lead to a grid-locked pattern (Fig. 2 and Movie S4, right).
When a growing open-ended wall encounters a branch point ahead of it, the outflowing stream at the end point of that wall is diverted symmetrically in perpendicular directions so that material can join the streams of the two walls that split away from the branch point (see Fig. 1, around the black circles). In effect, this diversion creates a local stagnation point: flows arrive from the direction of the open-ended wall and depart in the perpendicular direction, which is also the direction to which rods tend to align by flow-alignment. For bend walls—generated by an extensile stress—the director along the incident wall is perpendicular to the wall itself, and thus naturally aligned with the extensional axis of the flow at the stagnation point. We therefore propose that the structure around the stagnation point acts as a trap, and its stability is central to understanding how flow alignment promotes the dynamical arrest of active turbulence.
@@
Figure 3: Domain wall nodes and their pseudo-topology. In all diagrams, black directed lines represent nematic bending walls, with gray lines tracing the director and cyan indicating the bending energy. 33, Nodes are either startpoints (green, negative pseudo-defects), branchpoints (blue, negative pseudo-defects), or endpoints (red, positive pseudo-defects). The orange curves illustrate the paths excluding walls used to define the pseudo-charge (Appendix C). 33, Two ways in which a startpoint-endpoint pair of opposite charge can be born. The inverse processes of (3) (complete wall dissolution) and (3) (local wall completion) result in the annihilation of such pairs. 3, Wall branching gives birth to a branchpoint-endpoint pair, also of opposite charge. A connected pair may annihilate via branch retraction. 3, A branchpoint (3) transitions into a startpoint when one outgoing wall disjoins. The inverse process corresponds to a startpoint (3) joining with a bare wall. 3, Connected branchpoints can shrink their connecting branch, transiently creating a 2Q-2Q pseudo-charged structure, exchange outgoing walls and extend again in the perpendicular direction. The processes (33) are further illustrated in Animations S1 to S5.
IV Pseudo-topology of domain wall networks
IV.1 Wall nodes define pseudo-defects